Zinc occurs is sulfide ores. How many liters of SO2(g) are produced at STP for each kilogram of roasted ore that is 80% zinc sulfide?
Answer #1

Given that 80% of 1 kg ore is ZnS

Thus, Mass of ZnS = 800 g

Molar mass of ZnS = 97.474 g/mol

Moles of ZnS = 800 g*(1 mol / 97.474 g) = 8.21 mol

Reaction is:

2ZnS(s) + 3O2 (g) ---> 2ZnO(s) + 2SO2(g)

Thus, moles of SO2 produced(n) = moles of ZnS = 8.21 mol

At STP, Pressure (P) = 1 atm

Temperature (T) = 273.15 K

Gas constant (R) = 0.0821 L. atm / mol.K

From Ideal gas equation

PV = nRT

V = nRT/P = (8.21 mol * 0.0821 L. atm / mol.K * 273.15 K) / 1 atm = 184.05 L